If (1+x+x 2 ) 6n =a 0 +a 1 x+....then (a 0 +a 3 +a 6 +.....)= (a)3 6n-1 (B)3 6n+1 (c)3 6n -1 (D)3 6n+1 -1

If (1+x+x2)6n=a0+a1x+....then


(a0+a3+a6+.....)=


(a)36n-1


(B)36n+1


(c)36n-1


(D)36n+1-1

Grade:12

1 Answers

askiitian expert akshay singh
14 years ago

HI VAIBHAV,

You can do this problem using a short trick,which is used generally to solve objective problems involving n,this will save ur time and u'll get the correct answer.Do onething put n=1/3,so expression becomes (1+x+x2)2 =1+2x+3x2 +2x3+x4 now a0+a3=1+2=3...........now solve all the answers for n=1/3 u'll find that option (a)36n-1=3 is the correct answer.